OpenStudy (anonymous):

a sector of a circle has created from a central angle with the measure of 60 degrees if the diameter of a circle 6 inches what is the area of a sector?

OpenStudy (anonymous):

I don't understand?

OpenStudy (jdoe0001):

http://0.tqn.com/d/math/1/0/G/H/sector.gif

OpenStudy (anonymous):

So what's the answer?

OpenStudy (jdoe0001):

the answer is use the given values and plug them in the formula :)

OpenStudy (anonymous):

I don't know how to plug it in?

OpenStudy (jdoe0001):

the formula uses a "central angle" do you have one in the question?

OpenStudy (anonymous):

60? I think? The answer choses are 1.5 in^2 B.2in^2 C.6in^2 D.8in^2

OpenStudy (anonymous):

And they all have the unit of pie with it

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Diameter = 6 in., so radius = 3 in. Area = (60/360) * pi * (3 in.)^2 = (1/6) * pi * 9 in.^2 = 1.5(pi) in.^2

OpenStudy (anonymous):

Thanks!!(:

OpenStudy (mathstudent55):

You're welcome.